# Curiosity-Driven Co-Development of Action and Language in Robots Through Self-Exploration

### v2

@@ 17 Hourglass
Table S5: English words and their indexes in one-hot vectors.
The robot’s forward model’s encoding of these two voices has two parts. The first part of the encoding is an embedding and recurrent neural network. This part is identical for the command voice and the feedback voices, ensuring that tokens are interpreted consistently across sources. Note that this RNN is “nested” within the forward model’s RNN, such that each of the robot’s steps includes three steps of interpreting voices. See figure S3. In the second part of the encoding, outputs from the first part of the encoding are processed with unique linear layers to produce separate estimated posteriors. To generate a prediction of the next voices, $h^{q}_{t}$ and $a^{enc}_{t}$ are concatenated and decoded using two separate recurrent neural networks for the command voice and feedback voice. See details in table S6.
@@
Figure S3: Recurrent step shared by command voice and feedback voice.
@@
We measured the success-rates of these three types of robots in the six types of actions. The plots in figure 3 show the rolling average of success-rates of the three types robots from the beginning of training to the end of training after 60000 epochs, with 99% confidence intervals. Specifically, the plots show results of the robots regarding the goals with combinations of action, color, and shape which the robots were not shown in training, testing for the ability to generalize vocabulary and syntax to unlearned combinations.
@@
Experiment 2 tests the relationship between success-rates with learned goals and unlearned goals, specifically by robots using all curiosity. See figure 4. The left column shows success-rate plots of robots with learned actions, while the right column shows success-rate plots of robots with unlearned actions. The first row shows results for robots using the complete vocabulary: 6 actions, 6 colors, and 5 shapes. The second and third row show results for robots trained with smaller vocabularies. In each of the three situations, the robots are trained with one third of the possible goals, and tested with the other two thirds.
@@
As we predicted in hypothesis iiiiii, the robot’s success-rates with learned actions initiates earlier than its success-rates with unlearned actions. This suggests pairing sentences of words precedes generalization with compositionality. As we predicted in hypothesis $iv$, larger vocabularies lead to faster generalization. All three collections of robots had success-rates of approximately 100% with learned actions. Robots which were trained with 60 of the 180 possible goals with 6 actions, 6 colors, and 5 shapes had success-rates of approximately 90% with unlearned tasks. Robot which were trained with 25 of the 75 possible goals with 5 actions, 5 colors, and 3 shapes had success-rates of approximately 50% with unlearned tasks. And robots which were trained with 16 of the 48 possible goals with 4 actions, 4 colors, and 3 shapes has success-rates of approximately 30% with unlearned tasks. The ability to generalize quickly is enhanced with the size of the vocabulary in use.

### v3

@@ 17 Hourglass
Table S5: English words and indexes. The English words used and their positions in one-hot vectors.
The robot’s forward model’s encoding of these two voices has two parts. The first part of the encoding is an embedding and recurrent neural network. This part is identical for the command voice and the feedback voices, ensuring that tokens are interpreted consistently across sources. Note that this RNN is “nested” within the forward model’s RNN, such that each of the robot’s steps includes three steps of interpreting voices. See Fig. S4. In the second part of the encoding, outputs from the first part of the encoding are processed with unique linear layers to produce separate estimated posteriors. To generate a prediction of the next voices, $h^{q}_{t}$ and $a^{enc}_{t}$ are concatenated and decoded using two separate recurrent neural networks for the command voice and feedback voice. See details in table S6.
@@
Fig. S4: Recurrent step shared by command voice and feedback voice.
@@
We measured the success-rates of these three types of robots in the six types of actions. The plots in Fig. 2 show the rolling average of success-rates of the three types robots from the beginning of training to the end of training after 60000 epochs, with 99% confidence intervals. Specifically, the plots show results of the robots regarding the goals with combinations of verb, adjective, and noun which the robots were not shown in training, testing for the ability to generalize vocabulary and syntax to unlearned combinations.
@@
Experiment 2 tests the relationship between success-rates with learned goals and unlearned goals, specifically by robots using all curiosity. See Fig. 6. The left column shows success-rate plots of robots with learned actions, while the right column shows success-rate plots of robots with unlearned actions. The first row shows results for robots using the complete vocabulary: 6 verbs, 6 adjectives, and 5 nouns. The second and third row show results for robots trained with smaller vocabularies. In each of the three situations, the robots are trained with one third of the possible goals, and tested with the other two thirds.
@@
As we predicted in hypothesis iiiiii, the robot’s success-rates with learned actions initiates earlier than its success-rates with unlearned actions. This suggests pairing sentences of words precedes generalization with compositionality. As we predicted in hypothesis $iv$, larger vocabularies lead to faster generalization. All three collections of robots had success-rates of approximately 100% with learned actions. Robots which were trained with 60 of the 180 possible goals with 6 verbs, 6 adjectives, and 5 nouns had success-rates of approximately 90% with unlearned actions. Robot which were trained with 25 of the 75 possible goals with 5 verbs, 5 adjectives, and 3 nouns had success-rates of approximately 50% with unlearned actions. And robots which were trained with 16 of the 48 possible goals with 4 verbs, 4 adjectives, and 3 nouns has success-rates of approximately 30% with unlearned actions. The ability to generalize quickly is enhanced with the size of the vocabulary in use.
Movie S1. Example of training. Compares robot with curiosity mid-training and after training.
Statistical Analysis of U-Shaped Patterns
To quantify U-shaped learning in exception handling, we scored the U-shaped structure of success-rate’s trajectories identifying non-monotonic developmental patterns consistent with representational redescription (?). The method combines robust smoothing, normalized scaling, and piecewise isotonic regression to fit a two-phase model with a central valley.
Consider one robot’s rolling-average success rate over training epochs for goals which are exceptions. The U-shape score is computed as follows:
  1. 1.
    Burn-in removal. The first 10% of training data is removed to avoid initialization noise.
  2. 2.
    Smoothing. The curve is smoothed using a Savitzky–Golay filter with a window length of approximately 3% of the series, reducing spurious local fluctuations.
  3. 3.
    Normalization. The smoothed curve is linearly scaled to the $[0,1]$ range using the 5th and 95th percentiles to ensure robustness across success-rate ranges.
  4. 4.
    Valley localization. The minimum point $i_{M}$ is located between 20% and 80% of the sequence length.
  5. 5.
    Piecewise isotonic regression. For each candidate split point $k$ near the valley (within $\pm 25\%$ of the series), the left segment is fit with a decreasing isotonic regression and the right segment with an increasing isotonic regression. A cost function is minimized:
    $\text{Cost}(k)=\text{MSE}(k)+\lambda\cdot(\text{drift from valley})^{2}\cdot\text{MSE}_{\text{base}},$
    where $\lambda=2.0$ penalizes drifting too far from the identified valley. Indices of $k$ define the left peak $i_{L}$ and right peak $i_{R}.$
  6. 6.
    Score calculation. If the best split passes depth and width criteria (minimum 3% depth, 6% width), a composite U-score is computed:
    $\text{U-score}=0.6\cdot\text{improvement}+0.25\cdot\text{depth}+0.15\cdot\text{width},$
    where:
    • Improvement is the fractional MSE reduction relative to the best monotonic baseline fit.
    • Depth is the drop from the valley to the 90th percentile of the surrounding peaks.
    • Width is the relative proportion of the sequence before/after the valley.
  7. 7.
    Index reporting. Indices of the left peak $i_{L}$, valley $i_{M}$, and right peak $i_{R}$ are marked with red vertical lines in Figure 7.
To compare robots training with exceptions and without exceptions, we computed U-shape scores for each robot individually and compared the two groups using a one-tailed Welch’s $t$-test (unequal variances):
$t=\frac{\bar{x}_{1}-\bar{x}_{2}}{\sqrt{\frac{s_{1}^{2}}{n_{1}}+\frac{s_{2}^{2}}{n_{2}}}}.$
The resulting test statistic confirmed that robots trained with exceptions exhibited significantly stronger U-shaped profiles than those without, with $p=0.0025$.